OpenStudy (anonymous):

Find dy/dx if x^4+2x^2y^3+y^2=21

OpenStudy (jhannybean):

Take the derivative of the function, but wherever you see a y, multiply it by \(\frac{dy}{dx}\) or \(y'\).

OpenStudy (jhannybean):

\[\frac{dy}{dx}(x^4+2x^2y^3+y^2=21)\]Focus on the \(2x^2y^3\). Use the product rule here, which states \(f'g+g'f\), \(f(x) = 2x^2\) and \(g(x) = y^3\).

OpenStudy (anonymous):

Quick question: For 2x^2y^3, can we separate like this, 2(x^2y^3)?

OpenStudy (jhannybean):

No.

OpenStudy (jhannybean):

Two separate functions.

OpenStudy (anonymous):

Why?

OpenStudy (jhannybean):

Because the 2 is multiplying the x, and not the y. if it were multiplying both it would read as \(2(x^2y^3)\) in which we would use the product rule first, then multiply in the 2 after.

OpenStudy (anonymous):

Oh okay. That's the reason why I got the question wrong. I got it now. Thank you :) @Jhannybean

OpenStudy (jhannybean):

You figured out the rest then?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

Awesome \(\checkmark\)

OpenStudy (kainui):

Well I hate to be _that guy_ but you can perfectly fine separate it like that 2(x^2y^3) \[\Large \frac{d}{dx}(2*(x^2y^3) )=\frac{d}{dx}(2)*(x^2y^3)+2*\frac{d}{dx}(x^2y^3)\] Now we know that the derivative of 2 with respect to x is 0 cause it's a constant so that first term disappears and we're left with \[\Large 2\frac{d}{dx}(x^2y^3)\] and you can continue to do the product rule on this term. So now you can see why constants can come out of the derivative. =)

OpenStudy (anonymous):

If you don't mind, can you show me your work because I did that way and I kept getting the wrong answer...